# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ Appendix A Proof of Corollary 4
Let $\omega=\zeta\lambda$. It holds that $\zeta=\frac{\omega^{2}}{e^{-\omega}-1}$ by Eq. (22). Let us define $g(\omega)=\frac{\omega^{2}}{e^{-\omega}-1}$. Therefore, the derivative of $g(\omega)$ with respect to $\omega$ is given by $g^{\prime}(\omega)=\frac{2\omega(e^{-\omega}-1)+\omega^{2}e^{-\omega}}{(e^{-\omega}-1)^{2}}.$ Therefore, $g(\omega)$ is monotonically increasing in $(-\infty,\varpi)$ and monotonically decreasing in $(\varpi,\infty)$ in which $g^{\prime}(\varpi)=0$. Further, it holds that $(\varpi+2)e^{-(\varpi+2)}=2e^{-2}$. Hence, we have $\varpi=-W_{0}\left(-\frac{2}{e^{2}}\right)-2$. Therefore, the characteristics equation Eq. (22) has two distinct negative real roots for $\zeta\in(0,\frac{\varpi^{2}}{e^{-\varpi}-1})$. Further, suppose the solution for Eq. (21) is pure imaginary when $\omega=\gamma$, i.e., $\lambda=i\theta$. In this case, it holds that $\sin(-\gamma\theta)=0$ and $\cos(-\gamma\theta)=-\gamma\theta^{2}+1.$ Therefore, we have $\gamma=\frac{\pi^{2}}{2}$ and $\theta=\pm\frac{2}{\pi}$. By this means, we arrive at Corollary 4.
Appendix B Proof of Proposition 7
@@
By modeling the communication channel with a fixed service rate in pure asynchronous SGD as an $M/D/1$ queue, the convergence rate of the asynchronous SGD algorithm is given by $\lambda_{\text{Q}}\approx\frac{W_{0}{(-\mathbb{E}\{\breve{\tau}_{\text{Q}}\}\eta v)}}{\mathbb{E}\{\breve{\tau}_{\text{Q}}\}\eta}$, in which $\mathbb{E}\{\breve{\tau}_{\text{Q}}\}$ is given by Eq. (32). Therefore, the optimum number of workers in asynchronous SGD through a queue can be determined by solving the following optimization problem $\min_{K\leq K_{\text{max}}}~{}~{}~{}\frac{\text{Real}\left\{W_{0}\left(-K\eta vz(K)\right)\right\}}{\eta z(K)},$ in which $\check{K}=\frac{\mu_{Q}}{\nu}$ is the capacity of the queue and $z(K)=1+\frac{1}{2\check{K}}\left(\frac{K}{\check{K}-K}+2\right)$. Next, the optimal number of workers in this setting, denoted by $K^{*}_{Q}$, should satisfy that $0<K^{*}_{Q}\eta vz(K^{*}_{Q})\leq\frac{1}{e}$ such that the $W_{0}({-K^{*}_{Q}\eta vz(K^{*}_{Q})})$ is real since $z(K)$ is monotonically increasing with $K$. Therefore, we could focus on the following optimization problem:

### v2

@@ Appendix C Proof for Corollary 6
Therefore, $g(\omega)$ is monotonically increasing in $(-\infty,\varpi)$ and monotonically decreasing in $(\varpi,\infty)$ in which $g^{\prime}(\varpi)=0$. Further, it holds that $(\varpi+2)e^{-(\varpi+2)}=2e^{-2}$. Hence, we have $\varpi=-W_{0}\left(-\frac{2}{e^{2}}\right)-2$. Therefore, the characteristics equation Eq. (41) has two distinct negative real roots for $\zeta\in(0,\frac{\varpi^{2}}{e^{-\varpi}-1})$. Further, suppose the solution for Eq. (40) is pure imaginary when $\omega=\gamma$, i.e., $\lambda=i\theta$. In this case, it holds that $\sin(-\gamma\theta)=0$ and $\cos(-\gamma\theta)=-\gamma\theta^{2}+1.$ Therefore, we have $\gamma=\frac{\pi^{2}}{2}$ and $\theta=\pm\frac{2}{\pi}$. By this means, we arrive at Corollary 6.
Appendix D Proof of Proposition 12
Given the learning rate $\eta$, the two roots of the characteristics function is given by determined by Proposition 2. Further, it holds that $\text{Re}\{\lambda_{0}\}>\text{Re}\{\lambda_{1}\}$ when $\breve{\tau}\eta v<\frac{1}{e}$, and $\text{Re}\{\lambda_{0}\}=\text{Re}\{\lambda_{1}\}$ when $\breve{\tau}\eta v\geq\frac{1}{e}$. In this way, the convergence rate of the SGD algorithm is mainly determined by the dominant root $\text{Re}\{\lambda_{0}\}$, i.e., $\mathbb{E}\{\hat{x}(t)\}=A_{0}e^{\lambda_{0}{K\nu t}}+A_{1}e^{\lambda_{1}{K\nu t}}$ where $A_{0}$ and $A_{1}$ are constants while $K\nu$ corresponds to the rate of the gradient arrival process. Next, we adopt dominant pole approximation by focusing on the dominant pole $K\nu\lambda_{0}$, which has the largest effect on the system’s response. In particular, it holds that $\lambda_{0}=\frac{W_{0}{(-(K-1){v}\eta)}}{(K-1)\eta}.$ Such, $K^{*}=\arg\min_{K}\frac{K\text{Re}\{W_{0}(-(K-1){v}\eta)\}}{(K-1)\eta}$. Next, it holds that $\frac{K}{K-1}=1+\mathcal{O}(1/K)$ under the condition that $K\gg 1$. Notably, $\text{Re}\{W_{0}(s)\}$ attains its minimum at $s=-\frac{1}{e}$, leading us to the proposition.
Appendix E Proof of Proposition 14
@@
By modeling the communication channel with a fixed service rate in pure ASGD as an $M/D/1$ queue, the dominant root of the ASGD algorithm is given by $\lambda_{\text{Q}}=\frac{W_{0}{(-\mathbb{E}\{\breve{\tau}_{\text{Q}}\}\eta v)}}{\mathbb{E}\{\breve{\tau}_{\text{Q}}\}\eta}$, in which $\mathbb{E}\{\breve{\tau}_{\text{Q}}\}$ is given by Eq. (51). Therefore, the optimum number of workers in ASGD through a queue can be determined by solving the following optimization problem $\min_{K\leq K_{\text{max}}}~{}~{}~{}\frac{\text{Re}\left\{W_{0}\left(-K\eta vz(K)\right)\right\}}{\eta z(K)},$ in which $\check{K}=\frac{\mu_{Q}}{\nu}$ is the capacity of the queue and $z(K)=1+\frac{1}{2\check{K}}\left(\frac{K}{\check{K}-K}+2\right)$. Next, the optimal number of workers in this setting, denoted by $K^{*}_{Q}$, should satisfy that $0<K^{*}_{Q}\eta vz(K^{*}_{Q})\leq\frac{1}{e}$ such that the $W_{0}({-K^{*}_{Q}\eta vz(K^{*}_{Q})})$ is real since $z(K)$ is monotonically increasing with $K$. Therefore, we could focus on the following optimization problem: